A random sample of 100 students who took a statistics exam were asked their score on the exam. The mean score on the test was 50; the standard deviation was 10. The scores are normally distributed.
A) What percent of the students scored below 30?
B) What percent scored between 30 and 55?

Answer by stanbon(57356)

A random sample of 100 students who took a statistics exam were asked their score on the exam. The mean score on the test was 50; the standard deviation was 10. The scores are normally distributed.
A) What percent of the students scored below 30?
Find the t-score 30
t(30) = (30-50)/10 = -2
Find Prob(t<-2); you get 0.0241..
----------------------------
B) What percent scored between 30 and 55?
t(55) = (55-30) / 10 = 2.5
t(30) = (30-30) / 10 = 0
P(0 < t < 2.5) = 0.4929..
========================
Cheers,
Stan H.